OpenStudy (anonymous):

Convert 8mm=______cm Pleasee help :)))

OpenStudy (anonymous):

0.8

OpenStudy (unklerhaukus):

10 mm =1 cm

OpenStudy (unklerhaukus):

1 mm = 1/10 cm 8 mm = 8/10 cm

OpenStudy (anonymous):